# Adaptive Gradient Normalization and Independent Sampling for (Stochastic) Generalized-Smooth Optimization

### v2

@@ Abstract
Recent studies have shown that many nonconvex machine learning problems satisfy a generalized-smooth condition that extends beyond traditional smooth nonconvex optimization. However, the existing algorithms are not fully adapted to such generalized-smooth nonconvex geometry and encounter significant technical limitations on their convergence analysis. In this work, we first analyze the convergence of adaptively normalized gradient descent under function geometries characterized by generalized-smoothness and generalized PŁ condition, revealing the advantage of adaptive gradient normalization. Our results provide theoretical insights into adaptive normalization across various scenarios. For stochastic generalized-smooth nonconvex optimization, we propose Independent-Adaptively Normalized Stochastic Gradient Descent, which leverages adaptive gradient normalization, independent sampling, and gradient clipping to achieve an $\mathcal{O}(\epsilon^{-4})$ sample complexity under relaxed noise assumptions. Experiments on large-scale nonconvex generalized-smooth problems demonstrate the fast convergence of our algorithm.
@@ 1 Introduction
In modern machine learning, the convergence of gradient-based optimization algorithms has been well studied in the standard smooth nonconvex setting. However, it has been shown recently that the standard $L$-smoothness fails to hold in many nonconvex machine learning problems, including distributionally-robust optimization (DRO)(Levy et al., 2020; Jin et al., 2021), meta-learning (Nichol et al., 2018; Chayti & Jaggi, 2024) and language models (Liu et al., 2023; Zhang et al., 2019). Instead, these problems were shown to satisfy a so-called generalized-smooth condition, in which the smoothness parameter can scale with the gradient norm in the optimization process (Zhang et al., 2019).
@@
In the existing literature, various works have proposed different algorithms for solving generalized-smooth nonconvex optimization problems. Specifically, various works have demonstrated that deterministic first-order algorithms such as gradient descent, normalized gradient descent and clipped gradient descent can achieve $\mathcal{O}(\epsilon^{-2})$ iteration complexity under mild assumptions (Li et al., 2024; Zhang et al., 2019; Chen et al., 2023; Gorbunov et al., 2024; Vankov et al., 2024b; Reisizadeh et al., 2023). These complexity results match the lower bound obtained by classical first-order methods. In particular, Chen et al. (2023) empirically demonstrated that proper usage of adaptive gradient normalization can substantially accelerate convergence. However, the formal theoretical justification and understanding of adaptive gradient normalization is still lacking for first-order algorithms in generalized-smooth optimization.
@@
On the other hand, some other works studied first-order stochastic algorithms in generalized-smooth nonconvex optimization (Li et al., 2024; Zhang et al., 2019; 2020a). Specifically, one line of work focused on the classic stochastic gradient descent (SGD) algorithm (Li et al., 2024). However, in the generalized-smooth setting, the convergence analysis of SGD either relies on adopting very large batch size or involves large constants (Li et al., 2024). Moreover, the empirical performance of SGD (Chen et al., 2023) is often unstable due to the ill-conditioned smoothness parameter when the gradient is large. Another line of work focused on clipped SGD, which leverages gradient normalization and clipping to handle the generalized-smooth geometry (Zhang et al., 2019; 2020a; Reisizadeh et al., 2023; Koloskova et al., 2023b). Although clipped SGD has demonstrated superior performance in solving large-scale problems, the existing theoretical analysis has several limitations. First, to establish convergence guarantees in the stochastic setting, existing studies often rely on strong assumptions, such as the stochastic approximation error being almost surely bounded or the use of extremely large batch sizes. Second, the existing designs of clipped SGD adopt the standard stochastic gradient normalization scheme, which is not fully adapted to the function geometry characterized by the generalized-smooth condition.
@@
  • Q1: In deterministic generalized-smooth optimization, how does adaptive gradient normalization affect the convergence rate of first-order algorithm, e.g., under Polyak-Łojasiewicz-type conditions ?
@@ 1.1 Our Contributions
To understand the advantage of using adaptive gradient normalization, we first study the convergence rate of adaptive normalized gradient descent (AN-GD) in deterministic generalized-smooth optimization under the generalized Polyak-Łojasiewicz (PŁ) condition over a board spectrum of gradient normalization parameters. Our results reveal the interplay among learning rate, gradient normalization parameter and function geometry parameter, and characterize their impact on the type of convergence rate. In particular, our results reveal the advantage of using adaptive gradient normalization and provide theoretical guidance on choosing proper gradient normalization parameter to improve convergence rate.
@@ 2 Related Work
Generalized-Smoothness. The concept of generalized-smoothness was introduced by Zhang et al. (2019) with the $(L_{0},L_{1})$-smooth condition, which allows a function to either have an affine-bounded hessian norm or be locally $L$-smooth within a specific region. This definition was extended by Chen et al. (2023), who proposed the $\mathcal{L}_{asym}^{*}(\alpha)$ and $\mathcal{L}_{sym}^{*}(\alpha)$ conditions, controlling gradient changes globally with both a constant term and a gradient-dependent term associated with power $\alpha$, thus applying more broadly. Later, Li et al. (2024) introduced $\ell$-smoothness, which use a non-decreasing sub-quadratic polynomial to control gradient differences. Mishkin et al. (2024) proposed directional smoothness, which preserves $L$-smoothness along specific directions.
@@
Algorithms for Generalized-Smooth Optimization. Motivated by achieving comparable lower bounds presented in Arjevani et al. (2023) under standard assumptions, algorithms for solving generalized-smooth problems can be categorized into two main series. The first series focus on gradient descent methods with constant learning rate. Li et al. (2024) proved that GD, SGD converge with $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ complexity under generalized-smoothness. To ensure convergence, Li et al. (2024) adopted assumptions such as gradient upper-bound and bounded variance assumption.
@@
Another series of work focus on adaptive methods. In the nonconvex deterministic settings, Zhang et al. (2019; 2020a); Gorbunov et al. (2024); Vankov et al. (2024b) showed that clipped GD can achieve an iteration complexity of $\mathcal{O}(\epsilon^{-2})$ under mild assumptions. Later, Chen et al. (2023) proposed $\beta$-GD that achieves $\mathcal{O}(\epsilon^{-2})$ iteration complexity. Specially, in the convex setting, Vankov et al. (2024b) analyzed clipped and normalized gradient descent under mild assumptions. Gorbunov et al. (2024) studied smoothed gradient clipping, gradient descent with Polyak step-size rule, and triangle method by varying the learning rates. Both works achieve the best-known $\mathcal{O}({\epsilon}^{-1})$ convergence rate for generalized-smooth convex optimization In the nonconvex stochastic setting, Zhang et al. (2020b) proved that clipped SGD achieves an $\mathcal{O}(\epsilon^{-4})$ sample complexity under the standard $L$-smoothness and bounded variance assumptions. Also, Zhang et al. (2019) established the same convergence rate for clipped SGD under generalized smoothness and almost surely bounded noise assumption. Moreover, Reisizadeh et al. (2023); Koloskova et al. (2023b) analyzed the convergence of clipped SGD under the bounded variance assumption. Notably, Koloskova et al. (2023b) pointed out that, with bounded variance, clipped SGD only converge to a small neighborhood whose radius depends on data variance. To address this issue, Reisizadeh et al. (2023) employed a large batch size $\mathcal{O}(\epsilon^{-2})$ to suppress the impact of the variance term and achieve $\mathcal{O}(\epsilon^{-4})$ complexity. We summarize these results under mild assumptions in Table 1. There are also studies exploring the performance of other stochastic algorithms under the generalized-smooth condition. Wang et al. (2023); Faw et al. (2023); Hong & Lin (2024) studied AdaGrad Duchi et al. (2011b) under generalized-smooth and affine variance assumption with different learning rate schemes. They all attains $\mathcal{\tilde{O}}(1/\sqrt{T})$ convergence rate under mild conditions. Xie et al. (2024a) studied trust-region methods convergence under generalized-smoothness. Several works also studied stochastic acceleration methods under the generalized-smoothness condition. Zhang et al. (2020a) proposed a general clipping framework by leveraging momentum clipping and they achieve $\mathcal{O}(\epsilon^{-4})$ sample complexity under almost sure bounded noise assumption; Jin et al. (2021) studied normalized SGD with momentum under parameter-dependent achieves $\mathcal{O}(\epsilon^{-4})$ sample complexity under generalized-smooth and affine variance assumption; Hübler et al. (2024) studied normalized SGD with momentum Cutkosky & Mehta (2020) associated with parameter-agnostic learning rates, which establishes $\mathcal{O}(\frac{1}{\epsilon^{4}}\log(\frac{1}{\epsilon}))$ convergence rate under bounded variance assumption. By adjusting batch size, Chen et al. (2023); Reisizadeh et al. (2023) demonstrated that the SPIDER algorithm (Fang et al., 2018) can reach the optimal $\mathcal{O}(\epsilon^{-3})$ sample complexity under affine variance assumption for sum-type functions. Furthermore, Zhang et al. (2024b); Wang et al. (2024a; b); Li et al. (2023) explored the convergence of RMSprop (Hinton et al., 2012) and Adam (Kingma, 2014) under different noise assumptions, including coordinate-wise affine variance, expected affine variance, bounded variance and sub-gaussian distribution. The acceleration of sign-SGD was explored by Jiang et al. (2024); Sun et al. (2023b), and its generalized-smooth variants were subsequently analyzed by Crawshaw et al. (2022); Sun et al. (2023a), where they both obtain $\mathcal{O}(\epsilon^{-4})$ under almost-sure bound and bounded variance assumptions, respectively.
@@
Koloskova et al. (2023a) GS var bound Clip-SGD $\Omega(1)$ $\mathcal{O}(\frac{1}{\sqrt{T}})\text{ to }\min\{\sigma,\frac{\sigma^{2}}{c}\}$
@@
Table 1: Comparison of existing clipping-based algorithms. Denote the stochastic gradient and clipping threshold as $\nabla f_{\xi}(w_{t})$ and $c$, respectively. Explanation of abbreviations: (GS) refers to the generalized-smoothness condition, and (LS) refers to the standard $L$-smoothness condition. (a.s. bound) means that for every sample, the stochastic gradient bias is almost surely bounded, i.e., $\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|\leq\tau_{2}$; (var bound) denotes the bounded variance condition, $\mathbb{E}_{\xi}[\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|^{2}]\leq\tau_{2}^{2}$; (a.s. affine) implies that for every sample, the stochastic gradient bias satisfies the affine bound $\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|\leq\tau_{1}\|\nabla F(w_{t})\|+\tau_{2}$ almost surely, for some $10\leq\tau_{1}<$ 1 and $\tau_{2}>0$. The term Clip-SGD refers to algorithms that use stochastic gradient clipping of the form $w_{t+1}=w_{t}-\gamma\min\{1,\frac{c}{\|\nabla f_{\xi}(w_{t})\|}\}\nabla f_{\xi}(w_{t})$; SGD refers to stochastic gradient descent; General-Clip denotes algorithms that incorporate momentum during clipping and updates $w_{t+1}=w_{t}-(\nu\min\{c,\frac{\gamma}{\|m_{t}\|}\}m_{t}+(1-\nu)\min\{c,\frac{\gamma}{\|\nabla f_{\xi}(w_{t})\|}\}\nabla f_{\xi}(w_{t}))$, where $m_{t+1}=(1-\beta)m_{t}+\beta\nabla f_{\xi}(w_{t})$. The label $\epsilon$-stationary indicates that the algorithm can converge to an arbitrary target error, while $\mathcal{O}(\frac{1}{\sqrt{T}})\text{ to }\min\{\sigma,\frac{\sigma^{2}}{c}\}$ implies $\mathcal{O}(\frac{1}{\sqrt{T}})$ convergence to a neighborhood with radius $\min\{\sigma,\frac{\sigma^{2}}{c}\}$.
@@
Machine Learning Applications. Generalized-smoothness has been studied under various machine learning framework. Levy et al. (2020); Jin et al. (2021) studied the dual formulation of regularized DRO problems, where the loss function objective satisfies generalized-smoothness. Chayti & Jaggi (2024) identified their meta-learning objective’s smoothness constant increases with the norm of the meta-gradient. Gong et al. (2024b); Hao et al. (2024); Gong et al. (2024a); Liu et al. (2022b) explored algorithms for bi-level optimization and federated learning within the context of generalized-smoothness. Zhang et al. (2024a) developed algorithms for multi-task learning problem where the objective is generalized-smooth. Xie et al. (2024b) studied online mirror descent when the objective is generalized-smooth. Xian et al. (2024) studied min-max optimization algorithms’ convergence behavior under generalized-smooth condition. There is a concurrent work (Vankov et al., 2024a) using independent sampling with clipped SGD framework to solve variation inequality problem (SVI). Based on this idea, they also propose stochastic Korpelevich method for clipped SGD. Under generalized-smooth condition, they proved almost-sure convergence in terms of distance to solution set tailored for solving stochastic SVI problems.

### v3

@@ Abstract
Recent studies have shown that many nonconvex machine learning problems satisfy a generalized-smooth condition that extends beyond traditional smooth nonconvex optimization. However, the existing algorithms are not fully adapted to such generalized-smooth nonconvex geometry and encounter significant technical limitations on their convergence analysis. In this work, we first analyze the convergence of adaptively normalized gradient descent under function geometries characterized by generalized-smoothness and generalized PŁ condition, revealing the advantage of adaptive gradient normalization. Our results provide theoretical insights into adaptive normalization across various scenarios. For stochastic generalized-smooth nonconvex optimization, we propose Independent-Adaptively Normalized Stochastic Gradient Descent algorithm, which leverages adaptive gradient normalization, independent sampling, and gradient clipping to achieve an $\mathcal{O}(\epsilon^{-4})$ sample complexity under relaxed noise assumptions. Experiments [^1] Code available at github.com/ynyang94/Gensmooth-IAN-SGD on large-scale nonconvex generalized-smooth problems demonstrate the fast convergence of our algorithm.
@@ 1 Introduction
In modern machine learning, the convergence of gradient-based optimization algorithms has been well studied in the standard smooth nonconvex setting. However, it has been shown recently that the standard $L$-smoothness fails to hold in many nonconvex machine learning problems, including distributionally-robust optimization (DRO)(Levy et al., 2020; Jin et al., 2021), meta-learning (Nichol et al., 2018; Chayti & Jaggi, 2024) and language models (Liu et al., 2023; Zhang et al., 2019). Instead, these problems were shown to satisfy a so-called generalized-smooth condition (Zhang et al., 2019), in which the smoothness parameter can scale with the gradient norm in the optimization process.
@@
In the existing literature, various works have proposed different algorithms for solving generalized-smooth nonconvex optimization problems. Specifically, Li et al. (2024); Zhang et al. (2019); Chen et al. (2023); Gorbunov et al. (2024); Vankov et al. (2024b); Reisizadeh et al. (2023) have demonstrated that deterministic first-order algorithms such as gradient descent, normalized gradient descent and clipped gradient descent can achieve $\mathcal{O}(\epsilon^{-2})$ iteration complexity under mild assumptions. These complexity results match the lower bound obtained by classical first-order methods (Arjevani et al., 2023). In particular, Chen et al. (2023) empirically demonstrated that proper usage of adaptive gradient normalization can substantially accelerate convergence. However, the formal theoretical justification and understanding of adaptive gradient normalization is still lacking for first-order algorithms in generalized-smooth optimization.
@@
On the other hand, some other works, including Li et al. (2024); Zhang et al. (2019; 2020a) studied first-order stochastic algorithms in generalized-smooth nonconvex optimization. Specifically, one line of works (Li et al., 2024) focused on the classic stochastic gradient descent (SGD) algorithm (Ghadimi & Lan, 2013). However, in generalized-smooth setting, the convergence analysis of SGD either relies on adopting very large batch size or involves large constants (Li et al., 2024). These theoretical bottlenecks may be a key factor limiting the practical performance of SGD. Empirical observations from Chen et al. (2023) indicate that SGD often converges slowly due to ill-conditioned smoothness parameters when the gradient norm is large. Another line of works (Zhang et al., 2019; 2020a; Reisizadeh et al., 2023; Koloskova et al., 2023) focused on clipped SGD, which leverages gradient normalization and clipping to handle the generalized-smooth geometry. Although clipped SGD has demonstrated superior performance in solving large-scale problems, the existing theoretical analysis has several limitations. First, to establish convergence guarantees in the stochastic setting, existing studies often rely on strong assumptions, such as the stochastic approximation error being almost surely bounded or the use of extremely large batch sizes. Second, the existing designs of clipped SGD adopt the standard stochastic gradient normalization scheme, which is not fully adapted to the function geometry characterized by the generalized-smooth condition.
@@
  • Q1: In deterministic generalized-smooth optimization, how does adaptive gradient normalization affect the convergence rate of first-order algorithm, e.g., under Polyak-Łojasiewicz type conditions?
@@ 1.1 Our Contributions
To understand the advantage of using adaptive gradient normalization, we first study the convergence rate of adaptive normalized gradient descent (AN-GD) in deterministic generalized-smooth optimization under the generalized Polyak-Łojasiewicz (PŁ) condition over a broad spectrum of gradient normalization parameters. Our results reveal the interplay among learning rate, gradient normalization parameter and function geometry parameter, and characterize their impact on the type of convergence rate. In particular, our results reveal the advantage of using adaptive gradient normalization and provide theoretical guidance on choosing proper gradient normalization parameter to improve convergence rate.
@@ 2 Related Work
Generalized-Smoothness. The concept of generalized-smoothness was first introduced by Zhang et al. (2019) with the $(L_{0},L_{1})$-smooth condition, which allows a function to either have an affine-bounded Hessian norm or be locally $L$-smooth within a specific region. This definition was extended by Chen et al. (2023), who proposed the $\mathcal{L}_{asym}^{*}(\alpha)$ and $\mathcal{L}_{sym}^{*}(\alpha)$ conditions, controlling gradient changes globally with both a constant term and a gradient-dependent term associated with power $\alpha$, thus applying more broadly. Later, Li et al. (2024) introduced $\ell$-smoothness, which use a non-decreasing sub-quadratic polynomial to control gradient differences. Mishkin et al. (2024) proposed directional smoothness, which preserves $L$-smoothness along specific directions.
@@
Algorithms for Generalized-Smooth Optimization. Motivated by achieving comparable lower bounds presented in Arjevani et al. (2023) under standard assumptions, algorithms for solving generalized-smooth problems can be categorized into two main series. The first series focuses on gradient descent methods with constant learning rate. Li et al. (2024) proved that GD, SGD converge with $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ complexity under generalized-smoothness. To ensure convergence, Li et al. (2024) adopted assumptions such as gradient upper-bound and bounded variance assumption.
@@
Another series of work focuses on adaptive methods. In the nonconvex deterministic settings, Zhang et al. (2019; 2020a); Gorbunov et al. (2024); Vankov et al. (2024b) showed that clipped GD can achieve an iteration complexity of $\mathcal{O}(\epsilon^{-2})$ under mild assumptions. Later, Chen et al. (2023) proposed $\beta$-GD that achieves $\mathcal{O}(\epsilon^{-2})$ iteration complexity. Specifically, in the convex setting, Vankov et al. (2024b) analyzed clipped and normalized gradient descent under mild assumptions. Gorbunov et al. (2024) studied smoothed gradient clipping, gradient descent with Polyak step-size rule, and triangle method by varying the learning rates. Both works achieve the best-known $\mathcal{O}({\epsilon}^{-1})$ convergence rate for generalized-smooth convex optimization. In the nonconvex stochastic setting, Zhang et al. (2020b) proved that clipped SGD achieves an $\mathcal{O}(\epsilon^{-4})$ sample complexity under the standard $L$-smoothness and bounded variance assumptions. Also, Zhang et al. (2019) established the same convergence rate for clipped SGD under generalized smoothness and almost surely bounded noise assumption. Later, Reisizadeh et al. (2023); Koloskova et al. (2023) analyzed the convergence of clipped SGD under the bounded variance assumption. Notably, Koloskova et al. (2023) pointed out that, with bounded variance assumption, clipped SGD converges to $\epsilon$-stationary point with sample complexity $\mathcal{O}(\epsilon^{-5})$ when using $\Omega(1)$ samples per iteration. To improve convergence sample complexity, Reisizadeh et al. (2023) employed a large batch size $\mathcal{O}(\epsilon^{-2})$ to suppress the impact of the variance term and achieve $\mathcal{O}(\epsilon^{-4})$ complexity. We summarize these results under mild assumptions in Table 1.
A number of works also investigated the performance of diverse stochastic algorithms—extending beyond SGD and clipped SGD under the generalized-smooth condition. Wang et al. (2023); Faw et al. (2023); Hong & Lin (2024) studied AdaGrad (Duchi et al., 2011b) under generalized-smooth and affine variance assumption with different learning rate schemes. They all attain $\mathcal{\tilde{O}}(1/\sqrt{T})$ convergence rate under mild conditions. Xie et al. (2024a) studied trust-region methods convergence under generalized-smoothness. For stochastic acceleration methods under the generalized-smoothness condition, Zhang et al. (2020a) proposed a general clipping framework by leveraging momentum clipping and they achieve $\mathcal{O}(\epsilon^{-4})$ sample complexity under almost sure bounded noise assumption; Jin et al. (2021) studied normalized SGD with momentum under parameter-dependent learning rates schemes, which achieves $\mathcal{O}(\epsilon^{-4})$ sample complexity under generalized-smooth and affine variance assumption; Hübler et al. (2024) studied normalized SGD with momentum (Cutkosky & Mehta, 2020) under parameter-agnostic learning rates schemes, which establishes $\tilde{\mathcal{O}}({\epsilon^{-4}})$ convergence rate under bounded variance assumption. Chen et al. (2023); Reisizadeh et al. (2023) demonstrated that the SPIDER algorithm (Fang et al., 2018) can reach the optimal $\mathcal{O}(\epsilon^{-3})$ sample complexity under affine variance assumption for sum-type functions by employing a large batch size. Furthermore, Zhang et al. (2024b); Wang et al. (2024a; b); Li et al. (2023) explored the convergence of RMSprop (Hinton et al., 2012) and Adam (Kingma, 2014) under different noise assumptions, including coordinate-wise affine variance, expected affine variance, bounded variance and sub-gaussian distribution. The acceleration of sign-SGD was explored by Jiang et al. (2024); Sun et al. (2023b), and its generalized-smooth variants were subsequently analyzed by Crawshaw et al. (2022); Sun et al. (2023a), where they both obtain $\mathcal{O}(\epsilon^{-4})$ under almost-sure bound and bounded variance assumptions, respectively.
@@
Koloskova et al. (2023) GS var bound Clip-SGD $\Omega(1)$ $\epsilon$-stationary $\mathcal{O}(\epsilon^{-5})$
@@
Table 1: Comparison of existing clipping-based algorithms. Denote the stochastic gradient and clipping threshold as $\nabla f_{\xi}(w_{t})$ and $c$, respectively. Explanation of abbreviations: (GS) refers to the generalized-smoothness condition, and (LS) refers to the standard $L$-smoothness condition. (a.s. bound) means that for every sample, the stochastic gradient bias is almost surely bounded, i.e., $\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|\leq\tau_{2}$; (var bound) denotes the bounded variance condition, $\mathbb{E}_{\xi}[\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|^{2}]\leq\tau_{2}^{2}$; (a.s. affine) implies that for every sample, the stochastic gradient bias satisfies the affine bound $\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|\leq\tau_{1}\|\nabla F(w_{t})\|+\tau_{2}$ almost surely, for some $10\leq\tau_{1}<$ 1 and $\tau_{2}>0$. The term Clip-SGD refers to algorithms that use stochastic gradient clipping of the form $w_{t+1}=w_{t}-\gamma\min\{1,\frac{c}{\|\nabla f_{\xi}(w_{t})\|}\}\nabla f_{\xi}(w_{t})$; SGD refers to stochastic gradient descent; General-Clip denotes algorithms that incorporate momentum during clipping and updates $w_{t+1}=w_{t}-(\nu\min\{c,\frac{\gamma}{\|m_{t}\|}\}m_{t}+(1-\nu)\min\{c,\frac{\gamma}{\|\nabla f_{\xi}(w_{t})\|}\}\nabla f_{\xi}(w_{t}))$, where $m_{t+1}=(1-\vartheta)m_{t}+\vartheta\nabla f_{\xi}(w_{t})$. The label $\epsilon$-stationary indicates that the algorithm can converge to an arbitrary target error. For Koloskova et al. (2023), we report their convergence result under learning rate $\gamma=\epsilon^{2}/\left(\tau_{2}^{2}(L_{0}+cL_{1})\right)$ with $c=\tau_{2}/\epsilon$, where their analysis guarantees convergence to an $\epsilon$-stationary point.
@@
Machine Learning Applications. Generalized-smoothness has been studied under various machine learning frameworks. Levy et al. (2020); Jin et al. (2021) studied the dual formulation of regularized DRO problems, where the loss function objective satisfies generalized-smoothness. Chayti & Jaggi (2024) identified their meta-learning objective’s smoothness constant increases with the norm of the meta-gradient. Gong et al. (2024b); Hao et al. (2024); Gong et al. (2024a); Liu et al. (2022b) explored algorithms for bi-level optimization and federated learning within the context of generalized-smoothness. Zhang et al. (2024a) developed algorithms for multi-task learning problem where the objective is generalized-smooth. Xie et al. (2024b) studied online mirror descent when the objective is generalized-smooth. Xian et al. (2024) studied min-max optimization algorithms’ convergence behavior under generalized-smooth condition. There is a concurrent work (Vankov et al., 2024a) using independent sampling with clipped SGD framework to solve variational inequality problem (SVI). Based on this idea, they also propose stochastic Korpelevich method for clipped SGD. Under generalized-smooth condition, they proved almost-sure convergence in terms of distance to solution set tailored for solving stochastic SVI problems.